1. Solve for x
a) (x-1) / (x+2) + (x+7) / (x-40) = 0

Can you please help me out? Thanks so much in advance:)
Can you please show the steps it would really help me understand:)

Answer by stanbon(75887)

(x-1) / (x+2) + (x+7) / (x-40) = 0
------
Multiply thru by (x+2)(x-40) to get:
(x-1)(x-40)+(x+7)(x+2) = 0
-------
x^2-4x+40 + x^2+9x+14 = 0
------------------------------
2x^2 + 5x + 54 = 0
------
Use the Quadratic Formula.
No Real Number solutions
============================
Cheers,
Stan H.